OpenStudy (anonymous):

The container that holds the water for the football team is 1/9 full. After pouring in 6 gallons of water, it is 1/3 full. How many gallons can the container hold?

OpenStudy (anonymous):

Um, o-o

OpenStudy (paxpolaris):

if x is the number of gallons it can hold ...

OpenStudy (paxpolaris):

in the beginnning there is \(\frac19x\) gallons of water ... then you add 6 gallons and end up with a. \(\frac13x\)

OpenStudy (paxpolaris):

\[\Large \frac x9+6=\frac x3\]

OpenStudy (anonymous):

is it 27?

OpenStudy (paxpolaris):

\[\Large \checkmark\]

OpenStudy (anonymous):

thank you!! :)